Disk VSN: HOB+0119/32000/ | Data volume at beginning: 32159.422 GB
2018.192.18:00:00 Experiment started fine (Lucas).
2018.192.20:00:22.63 Module changed over to HOB+0117/32000/ (start position 30676.4 GB)
First scan on module 192-2000 (Lucas)
Drove out to Mt P and pumped steel I mean helium. I initially thought pressure/temp were stable and/or falling but they took a turn upwards above 100K and so needed refilling. While pump was on waited for temp < 100K and once it was I turned pump off and drove back. During drove temperature has fallen to 80K (Lucas).
Also took out full module and inserted new 6000 GB module
/home/www/auscope/opswiki/data/pages/handover/rd1805.txt · Last modified: 2018/07/11 22:12 by Atifur R. Khan